### v1

# Sonic4D: Spatial Audio Generation for Immersive 4D Scene Exploration

Abstract
Recent advancements in 4D generation have demonstrated its remarkable capability in synthesizing photorealistic renderings of dynamic 3D scenes. However, despite achieving impressive visual performance, almost all existing methods overlook the generation of spatial audio aligned with the corresponding 4D scenes, posing a significant limitation to truly immersive audiovisual experiences. To mitigate this issue, we propose Sonic4D, a novel framework that enables spatial audio generation for immersive exploration of 4D scenes. Specifically, our method is composed of three stages: 1) To capture both the dynamic visual content and raw auditory information from a monocular video, we first employ pre-trained expert models to generate the 4D scene and its corresponding monaural audio. 2) Subsequently, to transform the monaural audio into spatial audio, we localize and track the sound sources within the 4D scene, where their 3D spatial coordinates at different timestamps are estimated via a pixel-level visual grounding strategy. 3) Based on the estimated sound source locations, we further synthesize plausible spatial audio that varies across different viewpoints and timestamps using physics-based simulation. Extensive experiments have demonstrated that our proposed method generates realistic spatial audio consistent with the synthesized 4D scene in a training-free manner, significantly enhancing the immersive experience for users. Generated audio and video examples are available at https://x-drunker.github.io/Sonic4D-project-page.
1 Introduction
Benefiting from large-scale data available soomro2012ucf101 ; schuhmann2021laion ; yu2023celebv and recent advancement in generative models ho2020denoising ; song2020denoising , 4D generation (i.e., dynamic 3D scene generation) jiang2024consistentd ; ren2023dreamgaussian4d ; zeng2024stag4d ; zhu2025ar4d has emerged as a promising direction due to its powerful capability in modeling complex spatiotemporal dynamics of real-world scenes. By enabling spatiotemporally consistent renderings from arbitrary camera viewpoints, 4D generation facilitates various downstream applications such as AR/VR li20244k4dgen ; fritsch20173d , robotics khalid20224d ; hann20204d , and autonomous driving wang2024occsora ; min2024driveworld .
However, while existing 4D generation methods hann20204d ; liu2025free4d ; zhang20244diffusion ; zeng2024trans4d achieve impressive visual results, they typically neglect the generation of spatial audio consistent with the corresponding 4D scene (i.e., audio that varies with the listener’s viewpoint and follows physical acoustic principles), limiting the overall immersive experience to a large extent.
To address this limitation, in this paper we propose Sonic4D, a novel framework that enables free-viewpoint rendering of both dynamic visual content and spatially consistent audio, thereby ensuring more immersive audiovisual exploration within the generated 4D scene. To this end, as shown in Fig. 1, we design a three-stage pipeline: 1) Dynamic Scene and Monaural Audio Generation. To empower novel-view visual renderings and provide essential spatial priors for spatial audio generation, we first leverage a pre-trained 4D generative model to synthesize the dynamic 3D scene from a monocular video. In parallel, a video-to-audio generative model is utilized to produce monaural audio that is semantically aligned with the generated 4D scene, serving as the raw acoustic input for subsequent spatial audio rendering. 2) 3D Sound-Source Localization and Tracking. To enable accurate physical simulation of dynamic sound propagation, we further estimate the sound source’s trajectory in 3D space (i.e., the 3D locations of the sound source at different timestamps). To achieve this goal, we first use a multimodal large language model (MLLM) to perform pixel-level visual grounding on the input video, identifying the 2D coordinates of the sound source in each frame. These 2D positions are then back-projected onto the dynamic point cloud reconstructed in the previous stage, yielding a sequence of 3D coordinates that represent the sound source’s trajectory. 3) Physics-Driven Spatial Audio Synthesis. Given the estimated sound source location and the generated 4D environment, a physics-based simulation using acoustic room impulse responses (RIRs) is employed to simulate plausible spatial audio, enabling more immersive experiences in complex 4D scenes.
Notably, our method adopts a modular architecture that facilitates the integration of state-of-the-art pre-trained expert models, ensuring that the framework remains extensible and future-proof, with performance continually improving as more advanced expert models become available.
The main contributions of this paper can be summarized as follows:
  • We propose Sonic4D, a novel framework that achieves spatial audio generation for immersive 4D scene exploration. To the best of our knowledge, this is the first work that introduces spatial audio into the context of 4D generation.
  • We propose a three-stage framework to achieve spatial audio generation: 1) Dynamic Scene and Monaural Audio Generation to extract semantically aligned visual and audio priors from a single video; 2) 3D Sound-Source Localization and Tracking to recover sound source’s 3D trajectory for precise acoustic simulation; 3) Physics-Driven Spatial Audio Synthesis to render dynamic, viewpoint-adaptive binaural audio via physics-based room impulse response simulation.
  • Extensive experiments have demonstrated that our framework can effectively generate spatial audio consistent with 4D visual content, enabling much more immersive and coherent audiovisual experiences.
2 Related Works
2.1 4D Generation
Early advances in 4D generation have built upon techniques such as Score Distillation Sampling (SDS) poole2022sds , which leverages generative priors of pre-trained diffusion models to optimize dynamic scene representations. For example, MAV3D singer2023text4d proposed a multi-stage framework that sequentially optimizes static geometry using text-to-image diffusion, followed by motion synthesis using text-to-video models. This paradigm laid the groundwork for methods such as Dream-in-4D zheng2024unified , which further integrated motion priors from pretrained video diffusion models into dynamic NeRF-based representations pumarola2021d ; yan2023nerf . To improve rendering efficiency and motion fidelity, several methods have evolved from volumetric models to more expressive and compact formulations, such as AYG ling2024alignyourgaussians and Trans4D zeng2024trans4d , which employ deformation networks to animate static Gaussians across time. Concurrently, models like Consistent4D jiang2024consistentd and 4Diffusion zhang20244diffusion focused on enhancing spatiotemporal consistency, utilizing multiview interpolation, temporal alignment modules, or synchronized training strategies. Works such as Diffusion4D liang2024diffusion4d , PLA4D miao2024pla4d , and 4Dynamic yuan20244dynamic further emphasized controllability and efficiency by leveraging pixel-aligned supervision, video-based guidance, or hybrid representations combining mesh and Gaussian structures. Single-image driven methods like DreamGaussian4D ren2023dreamgaussian4d , Human4DiT shao2024human4dit , and Animate124 zhao2023animate124 extended monocular inputs into coherent 4D outputs using hierarchical architectures or diffusion-based motion refinement. Recent studies such as Free4D liu2025free4d and MVTokenFlow huang2025mvtokenflow further improve spatial-temporal coherence and visual quality without tuning large generative models.
Though achieving remarkable results, all these methods focus solely on visual content rendering, neglecting the generation of spatial audio that aligns with the dynamic scene. In contrast, in this paper we propose a novel framework for spatial audio generation that enables more coherent and immersive exploration of 4D scenes from arbitrary viewpoints.
2.2 Spatial Audio Generation
Early spatial audio generation methods  gao20192 ; xu2021visually ; leng2022binauralgrad predominantly took monaural inputs and “binauralized” them via neural networks conditioned on visual cues. In recent years, the advent of powerful generative architectures has spurred end-to-end models that accept text prompts sun2024both , images sun2024both ; dagli2024see , spatial parameters heydari2025immersediffusion ; kushwaha2025diff , silent videos kim2025visage , or full 360° panoramas liu2025omniaudio to jointly learn semantic and spatial audio features. For example, SpatialSonic sun2024both leverages spatial-aware encoders and azimuth state matrices within a latent diffusion framework to provide fine-grained multimodal spatial guidance for stereo audio generation. ViSAGe kim2025visage fuses CLIP visual embeddings with autoregressive neural audio codec modeling to generate coherent FOA directly from silent video frames. Recently, OmniAudio liu2025omniaudio addressed the 360° spatial audio gap by using a dual-branch architecture to fuse panoramic and FoV streams, enabling high-fidelity FOA generation from full spherical video content.
Although these methods benefit from the inherent capacity of generative models to capture complex multimodal distributions and yield impressive qualitative gains, they face two critical challenges. First, the data scarcity and domain gaps characteristic of end-to-end learning frameworks still constrain their ability to faithfully simulate spatial audio, resulting in suboptimal spatial rendering fidelity. Second, even when utilizing 360° video inputs, the synthesized sound field remains tied to the camera’s fixed pose and cannot account for dynamic listener motion or free-viewpoint changes within the scene.
Some alternative approaches chen2023novel ; liang2023av ; chen2024av tackle spatial audio by jointly reconstructing novel views and associated audio given ground-truth recordings at specific viewpoints; these methods, however, are restricted to static scenes and focus on view-specific audio re-rendering. In contrast, our work decouples semantic audio synthesis from spatial rendering and employs room impulse response (RIR) convolution to robustly simulate the spatial acoustics perceived by a moving listener in dynamic environments.
2.3 Sound Source Localization
Sound source localization aims to find the location of sound sources in an image or video frame. Initial studies senocak2018learning ; chen2021localizing ; senocak2023sound ; um2023audio addressed this problem by leveraging methods like cross-modal attention and contrastive learning to establish effective alignment between audio and visual modalities, mainly focusing on single sound source localization. Some works have taken this further to achieve multi-sound source localization hu2020discriminative ; mo2023audio ; kim2024learning , enabling the simultaneous localization of multiple sound sources from mixed audio and visual inputs. However, these approaches focus on exploring audio–visual alignments and cannot localize sound sources without ground-truth audio. Moreover, due to the complexity of three- and even four-dimensional scenarios, few works he2024sound3dvdet has addressed sound source detection or localization in 3D space.
Recently, the rapid advancement of Multimodal Large Language Models zhao2023bubogpt ; munasinghe2024videoglamm ; li2024groundinggpt (MLLMs) has made it possible to perform audio-visual grounding with strong generalization capabilities. In this work, we leverage the powerful prior knowledge of pretrained MLLMs to directly perform pixel-level visual grounding on the video input for sound source localization, followed by back-projection into 3D space to achieve accurate 3D sound source positioning and further spatial audio simulation.
3 Methods
Figure 1: Pipeline of Sonic4D. Our method is composed of three stages: 1) Dynamic Scene and Monaural Audio Generation: extracting semantically aligned visual scenes and audio priors from monocular videos (Sec. 3.1); 2) 3D Sound-Source Localization and Tracking: estimating the sound source’s trajectory in 3D space for physically accurate sound propagation modeling (Sec. 3.2); 3) Physics-Driven Spatial Audio Synthesis: leveraging a physics-based room impulse response simulation to realize spatial audio simulation. (Sec. 3.3)
Given a monocular input video $\bm{V}^{s}=\{\bm{I}^{s}_{i}\}_{i=1}^{n}\in\mathbb{R}^{n\times 3\times H\times W}$, our goal is to generate both novel-view videos and corresponding spatial audio given arbitrary target camera trajectory $\bm{T}^{r}=\{\bm{T}_{i}^{r}\}_{i=1}^{n}\in\mathbb{R}^{n\times 4\times 4}$, where each matrix $\bm{T}_{i}^{r}$ denotes the homogeneous camera pose at time $i$. As shown in Fig. 1, the overall pipeline is divided into three stages: 1) Dynamic scene and monaural audio generation; 2) 3D sound-source localization and tracking; 3) Physics-driven spatial audio synthesis. Each stage is discussed in detail in the following subsections.
3.1 Stage I: Dynamic Scene and Monaural Audio Generation
To obtain a novel-view video $\bm{V}^{r}$ while simultaneously capturing dynamic spatial priors for spatial audio simulation, a model that can produce high-fidelity, free-viewpoint renderings of dynamic scenes with strong spatiotemporal consistency is required. Motivated by recent progress in 4D generation, we adopt TrajectoryCrafter yu2025trajectorycrafter , a pretrained video generative model capable of synthesizing high-fidelity videos along arbitrary camera trajectories from a single-view input video, as our 4D content generator.
Specifically, we first apply a monocular depth estimator to estimate the depth maps $\bm{D}^{s}=\{\bm{D}^{s}_{i}\}_{i=1}^{n}\in\mathbb{R}^{n\times h\times w}$ of the input video $\bm{V}^{s}$. Subsequently, we lift $\bm{V}^{s}$ into a set of dynamic point clouds $\bm{P}=\{\bm{P}_{i}\}_{i=1}^{n}$ using inverse perspective projection $\bm{\Phi}^{-1}$, which is formulated as follows:
$\bm{P}_{i}=\Phi^{-1}([\bm{I}_{i}^{s},\bm{D}_{i}^{s}],\bm{K}),$ (1)
where $\bm{K}\in\mathbb{R}^{3\times 3}$ is the camera intrinsic matrix. Using the estimated dynamic point clouds, we synthesize novel-view renderings $\bm{I}^{r}=\{\bm{I}^{r}_{i}\}_{i=1}^{n}$ by projecting $\bm{P}$ onto the target camera trajectory $\bm{T}^{r}=\{\bm{T}^{r}_{i}\}_{i=1}^{n}\in\mathbb{R}^{n\times 4\times 4}$ using the following equation:
$\bm{I^{r}_{i}}=\Phi\left(\bm{T^{r}_{i}}\cdot\bm{P_{i}},\bm{K}\right),$ (2)
where $\bm{\Phi}$ denotes the projection operation. The rendered results $\bm{I}^{r}$ are then used as conditioning inputs to the video diffusion model of TrajectoryCrafter, guiding the generation of high-fidelity novel-view videos.
Next, we prepare the raw monaural audio signal required for subsequent physics-based spatial audio simulation. To this end, we adopt MMAudio cheng2024taming , a transformer-based video-to-audio synthesis model, to generate monaural audio $\bm{A}^{m}\in\mathbb{R}^{1\times T}$ that is temporally and semantically aligned with the input video $\bm{V}^{s}$. By decomposing the task of immersive 4D scene exploration into two sub-tasks—i.e., 4D scene generation and spatial audio synthesis—and leveraging pretrained expert models to extract the required visual and acoustic information, we can achieve greater flexibility and continually improve performance as upstream models evolve.
3.2 Stage II: 3D Sound Source Localization and Tracking
Figure 2: Illustration of Stage II: 3D Sound Source Localization and Tracking. We feed the video input frame-wisely into GroundingGPT li2024groundinggpt to obtain the sound source pixel coordinates. Then, we compute the average 3D coordinates of an $r\times r$ patch centered around the estimated 2D location via back-projection, and treat this as the 3D position of the sound source. Finally, we apply the DBSCAN ester1996density algorithm to filter and interpolate the sound-source trajectory, yielding a smooth and reliable final trajectory.
With the dynamic point cloud $\bm{P}$ and its associated monaural audio $\bm{A}^{m}$, we proceed to localize and track the sound source in 3D space to facilitate physics-based spatial audio generation. However, directly tracking the sound source in a 4D environment remains highly challenging, particularly in highly dynamic scenes with complex geometry. To address this, we first localize the sound source in the 2D pixel space and then back-project the estimated 2D coordinates into 3D space, thereby modeling the sound source trajectory within the generated dynamic scene.
To achieve this goal, existing 2D sound-source localization methods um2023audio ; senocak2023sound ; park2024can typically require ground-truth audio inputs and are often tailored to specific scenarios and input sizes. This dependence on “true” audio runs counter to Sonic4D’s very generation-centric design. In contrast, given the rich commonsense knowledge and reasoning abilities of multimodal large language models (MLLMs), we propose to leverage GroundingGPT li2024groundinggpt , a highly generalizable MLLM focused on grounding tasks and trained on a variety of audio–visual datasets for sound source localization, as our 2D sound source localization component.
Concretely, as shown in Fig. 1, we craft a textual prompt:
In this image, there is a possible sound source. Please find it and estimate its position. The only value you need to return is the position of the sound source in the format [x0, y0, x1, y1]. Do not include any text other than that.
and feed, frame by frame, the source video $\bm{V}^{s}$ together with the prompt into GroundingGPT li2024groundinggpt . Assuming the grounding result for the $i$-th frame $\bm{I}_{i}^{s}$ is represented by a bounding box $[x_{0},y_{0},x_{1},y_{1}]$, we define the sound source pixel coordinates as the center of this bounding box, i.e.:
$(u_{i},v_{i})=\biggl{(}\Bigl{\lfloor}\tfrac{x_{0}+x_{1}}{2}\times W\Bigr{\rfloor},\;\Bigl{\lfloor}\tfrac{y_{0}+y_{1}}{2}\times H\Bigr{\rfloor}\biggr{)}.$ (3)
where $W$ and $H$ are the width and height of $\bm{V}^{s}$. Let $d_{i}=\bm{D}(u_{i},v_{i})$ denote the depth at pixel $(u_{i},v_{i})$. Then the 3D point of the sound source in camera coordinate system corresponding to frame $i$ is given by
$\bm{X}_{i}\;=\;\pi^{-1}\bigl{(}(u_{i},v_{i}),\,d_{i}\bigr{)}.$ (4)
Here, $\pi^{-1}$ denotes the back-projection operator that maps a pixel coordinate and its depth into a 3D point. Considering the sporadic errors introduced by monocular depth estimation at individual pixels, we compute the 3D coordinates of an $r\times r$ pixel patch surrounding $(u_{i},v_{i})$ and average them to obtain the sound-source’s 3D location as shown in Figure 2:
$\bar{\bm{X}}_{i}\;=\;\frac{1}{\lvert\mathcal{P}^{r}(u_{i},v_{i})\rvert}\sum_{(u,v)\in\mathcal{P}^{r}(u_{i},v_{i})}\pi^{-1}\bigl{(}(u,v),\,\bm{D}(u,v)\bigr{)}.$ (5)
where $\mathcal{P}^{r}(u_{i},v_{i})$ denotes the set of all pixels $(u,v)$ such that $\lvert u-u_{i}\rvert\leq(r-1)/2$ and $\lvert v-v_{i}\rvert\leq(r-1)/2$. Nonetheless, due to uncertainty in the bounding-box predictions, the back-projected trajectory $\{\bar{\mathbf{X}}_{i}\}$ may still exhibit significant fluctuations. To mitigate this, we apply DBSCAN ester1996density (Density-Based Spatial Clustering of Applications with Noise), a clustering-based outlier detection algorithm that groups dense regions while marking sparse points as noise, to filter out spurious estimates. Points marked as noise are then temporally filled in via linear interpolation between the nearest non–noise neighbours to produce a smooth trajectory $\textbf{Traj}_{src}=\{\bm{X}_{i}\}_{i=1}^{n}$.
In addition to the sound source’s trajectory, the receiver trajectory is also required for acoustic simulation. We use the user-specified camera trajectory as the receiver trajectory. For each frame $i$, the camera pose is defined as
$\bm{T}^{r}_{i}=\begin{bmatrix}\bm{R}^{r}_{i}&\bm{t}^{r}_{i}\\[3.00003pt] \bm{0}^{\mathsf{T}}&1\end{bmatrix},$ (6)
where $\bm{R}^{r}_{i}\in\mathbb{R}^{3\times 3}$ is the rotation matrix and $\bm{t}^{r}_{i}\in\mathbb{R}^{3}$ is the translation vector. The receiver’s 3D position is therefore $\bm{r}_{i}\;=\;\bm{t}^{r}_{i},$and the full receiver trajectory is $\textbf{Traj}_{rsv}\;=\;\{\bm{r}_{i}\}_{i=1}^{n}.$
3.3 Stage III: Physics-Driven Spatial Audio Synthesis
Given the sound source’s trajectory $\textbf{Traj}_{src}$, the receiver trajectory $\textbf{Traj}_{rsv}$, and the monaural audio signal $\bm{A}^{m}$, we perform a physics-based spatial audio simulation.
A central element of our simulation is the computation of Room Impulse Responses (RIRs) via the Image Source Method (ISM) allen1979image . Building on the Image-Source Method for static RIRs, we apply gpuRIR diaz2021gpurir to simulate dynamic sources and receivers by segmenting the trajectories into short intervals over which they are assumed stationary. Concretely, let the total signal length be $N$ samples and split it into $M$ segments at sample indices
$0=n_{0}<n_{1}<\cdots<n_{M}=N,$
For each segment $i\in[1,M]$, let $\mathbf{x}^{s}_{i}=\textbf{Traj}_{src}[i]$ denote the position of the sound source and $\mathbf{x}^{r}_{i}=\textbf{Traj}_{rsv}[i]$ denote the position of the receiver. We simulate binaural microphone placement by offsetting from $\mathbf{x}^{r}_{i}$ in the direction parallel to the camera’s rendering plane, yielding the left and right microphone positions $\mathbf{x}^{r}_{i,L}$ and $\mathbf{x}^{r}_{i,R}$. We compute the room impulse responses at the left and right microphones using the Image Source Method (ISM):
$h_{i,L}(\tau)=\mathrm{ISM}\bigl{(}\mathbf{x}^{s}_{i},\;\mathbf{x}^{r}_{i,L}\bigr{)},\quad h_{i,R}(\tau)=\mathrm{ISM}\bigl{(}\mathbf{x}^{s}_{i},\;\mathbf{x}^{r}_{i,R}\bigr{)}.$ (7)
Let the mono audio block for segment $i$ be
$b_{i}[k]=\bm{A}^{m}[n_{i}+k],\quad 0\leq k<n_{i+1}-n_{i}.$ (8)
Then the left and right binaural signals are obtained by block-wise convolution:
$y_{L}[n]=\sum_{i=0}^{M-1}\bigl{(}b_{i}*h_{i,L}\bigr{)}[\,n-n_{i}\,],\qquad y_{R}[n]=\sum_{i=0}^{M-1}\bigl{(}b_{i}*h_{i,R}\bigr{)}[\,n-n_{i}\,],$ (9)
where $h_{i,L}$ and $h_{i,R}$ are the left- and right-ear impulse responses in $h_{i}$. Putting it all together, the final binaural signal is
$\bm{A}^{b}=\bigl{(}y_{L}[n],\,y_{R}[n]\bigr{)}.$ (10)
To ensure the output audio is in a valid range for playback, we normalize the signal to the range $[-1,1]$:
$\bm{A}^{b}\leftarrow\frac{\bm{A}^{b}}{\max\left(\left|\bm{A}^{b}\right|\right)}.$ (11)
Compared with learning-based methods for spatial audio generation, our physics-based audio rendering ensures physical plausibility by explicitly modeling sound propagation using room impulse responses, resulting in more realistic spatial cues such as directionality and reverberation.
By decomposing the task into three modular stages, our framework maximizes flexibility and extensibility. Each stage benefits from pretraining or physical priors, allowing Sonic4D to scale with future advances in generative modeling and spatial simulation. Moreover, this decoupled design enables plug-and-play improvements, making Sonic4D suitable for a broad range of content creation and immersive applications.
4 Experiments
4.1 Experimental Settings
Since this is the first attempt to perform spatial audio generation for immersive 4D scene exploration, existing evaluation benchmarks are not directly applicable. Therefore, we conducted a user preference study to assess the perceptual quality of spatial audio generated by our proposed and other methods. We choose MMAudio cheng2024taming as the baseline method, whose generated audio is monaural and contains no spatialization.
To determine whether our generated spatial audio conveys a sense of space, we paired the same viewpoint-specific video with either the spatialized stereo audio or the non-spatialized mono audio for comparison. Participants were first asked, for each video pair (identical visuals but different audio), to choose which audio track felt more spatial. Next, they rated both audio tracks on two criteria from 1 to 5:
  • Spatial Localization Accuracy: Measures how faithfully the interaural level and time-difference cues recreate the true 3D source position. Participants rate from 1 (“no sense of source direction”) to 5 (“precise and stable perception of source location”).
  • Audio–Visual Spatial Congruence: Assesses how consistently the evolving audio cues (e.g. changes in loudness or binaural disparity) track the on-screen motion of the subject or camera. Participants rate from 1 (“audio movement conflicts with visuals”) to 5 (“audio shifts perfectly match visual motion”).
We recruited 25 participants, each of whom evaluated 33 video-audio pairs. All participants completed the experiment in a quiet environment while wearing binaural headphones. To more comprehensively demonstrate our model’s capabilities, we categorized the videos based on custom camera trajectories into static and dynamic viewpoints. Static viewpoints indicate using the original camera view unchanged or only modest transformations of that original view, with the viewpoint remaining fixed throughout; Dynamic viewpoints include translational camera movements, arc-shaped camera trajectories, and push-in/pull-back motions.
For data processing, we first excluded participants whose ratings exhibited low Pearson correlation with the overall preference distribution. We then computed and reported the preference percentages and mean opinion scores (MOS) across all viewpoints, as well as separately for static and dynamic viewpoints.
4.2 Quantitative Comparisons
Metric All Static viewpoints Dynamic viewpoints
Preference (%)
   MMAudio cheng2024taming 10.97% 12.41% 9.90%
   Sonic4D (ours) 89.03% 87.59% 90.10%
MOS-SLA
   MMAudio cheng2024taming 2.322 2.357 2.296
   Sonic4D (ours) 4.013 3.986 4.033
MOS-AVSC
   MMAudio cheng2024taming 2.418 2.493 2.363
   Sonic4D (ours) 3.977 3.980 3.975
Table 1: Subjective evaluation results for Sonic4D vs. MMAudio cheng2024taming .
We present the results of user study evaluation comparing the spatial audio rendered by our method (Sonic4D) against the non-spatialized mono audio from MMAudio cheng2024taming . As shown in Table 1, Sonic4D achieved a strong overall preference rate of 89.03%, indicating that the spatial audio generated by Sonic4D exhibits a significant perceptual difference in spatiality compared to the original mono audio, thus demonstrating the effectiveness of our model. This trend held consistently across both static viewpoints (87.59%) and dynamic viewpoints (90.10%), demonstrating the robustness of our approach under varying camera trajectories.
In terms of Mean Opinion Scores (MOS), our method significantly outperforms the baseline in both evaluation dimensions. For Spatial Localization Accuracy (MOS-SLA), Sonic4D achieved a mean score of 4.013, compared to MMAudio cheng2024taming ’s 2.322. This suggests that participants were clearly able to perceive accurate and stable sound source positions when listening to our spatialized audio. A similar advantage is observed for Audio–Visual Spatial Congruence (MOS-AVSC), where Sonic4D scored 3.977 versus 2.418 for MMAudio cheng2024taming , indicating a much stronger alignment between auditory motion cues and on-screen visual dynamics.
Overall, these results validate that Sonic4D produces compelling and perceptually convincing spatial audio that generalizes well across diverse viewing conditions. Participants not only preferred the spatialized version, but also rated it higher in terms of spatial realism and audio-visual coherence, supporting the effectiveness of our approach in rendering viewpoint-consistent spatial audio.
Figure 3: Qualitative results across different scenarios. We present comparisons of the spatial audio generated by Sonic4D conditioning on various camera trajectories, including static camera viewpoints, camera circling around the subject, rightward panning, and pulling out. These examples demonstrate the temporal and spatial alignment between the generated spatial audio and the motion of the visual subject.
4.3 Qualitative Comparisons
As shown in Fig. 3, we present additional qualitative comparisons with the baseline method. We predefined three types of camera trajectories: static viewpoint (case 1), camera circling around the subject (case 2), rightward panning and pulling out (case 3). For each case, we show the rendered video from the novel viewpoint, annotated with the motion direction of the visual subject within the frame. We then present the original waveform along with the stereo spatial audio waveform generated by Sonic4D. We highly recommend using headphones to listen to the specific examples provided in our supplementary material.
In case 1, the helicopter first moves leftward and then rightward across the frame, resulting in a perceived spatial effect where the sound shifts from being centered, to the left, and then to the right. By comparing Regions 1 and 2 with Regions 3 and 4 in the spatial audio waveform, we observe that the left channel initially has greater amplitude than the right. As the helicopter moves rightward, the left channel amplitude rapidly decreases and eventually falls below the right, aligning well with the expected spatial perception of the moving object. In case 2, the camera starts on the right side of a ukulele player and orbits to the left. The left-right disparity in Regions 1 and 2 , as well as the temporal progression within those regions, closely mirrors the trajectory, indicating strong spatial tracking in the generated audio. Case 3 demonstrates that the same source video can yield notably different spatial audio depending on the camera motion, validating the view-dependence of our model. In video 1, as the camera pans rightward, the fountain becomes closer to the left ear, leading to increased amplitude in the left channel (Regions 1,2 ). In video 2, the camera gradually pulls out, producing a fading splash sound that matches the increasing distance (Regions 3,4,5,6 ).
These qualitative results demonstrate that the spatial audio generated by Sonic4D aligns closely with the visual cues, reflecting both spatial position and dynamic changes in the scene.
5 Conclusions
In this paper, we present Sonic4D, a novel paradigm that enables spatial audio generation for immersive 4D scene exploration. Specifically, our method consists of three stages: 1) Dynamic Scene and Monaural Audio Generation: we first employ pre-trained expert models to generate the 4D scene and its corresponding monaural audio, aiming to provide sufficient spatial and acoustic priors for subsequent spatial audio simulation. 2) 3D Sound-Source Localization and Tracking: we further propose a pixel-level visual grounding strategy to estimate the sound source’s trajectory in 3D space, facilitating spatial alignment between audio and visual content in dynamic scenes. 3) Physics-Driven Spatial Audio Synthesis: Based on the estimated sound source’s trajectory, we synthesize plausible, viewpoint-adaptive spatial audio using physics-based room impulse response (RIR) simulation. Extensive experiments demonstrate that our proposed method generates realistic spatial audio that is consistent with the synthesized 4D scene in a training-free manner, significantly enhancing the user’s immersive experience.
6 Limitations
Although our method can generate spatial audio that follows the camera viewpoint with clear directional cues, it still has several potential limitations:
  • While we model the dynamic trajectories of both the sound source and the listener using physics-based methods, the quality of the synthesized binaural audio may still be limited due to the lack of fine-grained acoustic simulation, such as incorporating Head-Related Transfer Functions (HRTF) or complex wave-based propagation models.
  • For simplicity, we constrain our setup to a single sound source per video. However, thanks to the flexibility of our framework, multi-source localization and generation can be supported by modifying the prompts provided to GroundingGPT and MMAudio accordingly.
  • Our use of monocular depth–derived point clouds lacks detailed geometric modeling, which may limit the physical accuracy of spatial audio propagation in complex environments.
Appendix A Details of User Study
A.1 Experiment Setup and Procedure
The user study consists of three main stages:
Step 1: Dynamic Scene and Monaural Audio Generation. In the first stage of our pipeline, as described in the Dynamic Scene and Monaural Audio Generation section, we used TrajectoryCrafter [54] to predict dynamic point clouds for 4D scene construction. Since TrajectoryCrafter [54] uses the relative transformation between the user-specified camera trajectory and the source camera as the basis for novel view rendering and requires a stabilized source video, we selected only source videos that meet this requirement. To generate these videos, we manually composed prompts for commercial text-to-video models (e.g., Kling [60], Dreamina [61], Vidu [62]), and for each sample, we designed a custom camera trajectory to render the target view. We then leveraged Sonic4D to generate both the novel view video and the corresponding spatialized audio.
Step 2: Interface Setup and Counterbalancing. As illustrated in Figure4, we created a local web-based interface to facilitate user evaluation. The landing page provided clear instructions on the task and rating criteria as shown in 4(a). We created 33 question pages for participants to complete. To ensure counterbalancing, we randomized the order of all audio-video pairs as shown in 4(b), as well as the internal ordering of each pair (left/right assignment), to mitigate any potential order effects or positional biases. Participants were asked to rate each video on the following two aspects using a 5-point Likert scale. The detailed rating criteria are as follows:
  • Spatial Localization Accuracy:
    • 5 – The position of the subject or camera can be clearly perceived from the audio.
    • 4 – The direction or positional changes can generally be perceived.
    • 3 – Some spatial positioning is conveyed through the audio.
    • 2 – Audio provides vague or weak spatial cues.
    • 1 – Little to no sense of spatial position from the audio.
  • Audio–Visual Spatial Congruence:
    • 5 – The direction indicated by audio cues (e.g., volume changes, binaural differences) perfectly matches the visual motion.
    • 4 – The audio directionality mostly matches the visuals.
    • 3 – The audio and visual alignment is partially consistent.
    • 2 – The audio and visuals are misaligned or direction is hard to perceive.
    • 1 – The spatial audio cues conflict with or fail to reflect the visuals.
(a) User study instruction page outlining participation guidelines and evaluation criteria.
(b) Example of a user study question page, displaying paired videos and corresponding rating options.
Figure 4: Screenshots of the user study interface.
Step 3: Data Collection and Filtering. After collecting all participant responses regarding spatial audio preference and ratings, we computed the Pearson correlation between each participant’s responses and the overall average responses. Specifically, for each participant $i$, we encode their answers on all preference questions as a numerical vector $\mathbf{a}^{(i)}$, where each answer is mapped to a scalar:
$a^{(i)}_{j}=\begin{cases}1,&\text{if the participant preferred Video 1}\\ 2,&\text{if the participant preferred Video 2}\\ 0,&\text{if the participant selected ``Almost same''}\end{cases}$ (12)
Here, $a^{(i)}_{j}$ denotes the response of the $i$-th participant to the $j$-th question, where $j=1,2,\dots,M$. The full response vector for participant $i$ is:
$\mathbf{a}^{(i)}=[a^{(i)}_{1},a^{(i)}_{2},\dots,a^{(i)}_{M}]$
We then compute the mean preference vector across all $N$ participants:
$\bar{\mathbf{a}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{a}^{(i)}$ (13)
Next, the Pearson correlation coefficient between each participant’s response vector and the mean vector is computed as:
$r^{(i)}=\frac{\sum_{j=1}^{M}\left(a^{(i)}_{j}-\bar{a}^{(i)}\right)\left(\bar{a}_{j}-\bar{\bar{a}}\right)}{\sqrt{\sum_{j=1}^{M}\left(a^{(i)}_{j}-\bar{a}^{(i)}\right)^{2}}\sqrt{\sum_{j=1}^{M}\left(\bar{a}_{j}-\bar{\bar{a}}\right)^{2}}}$ (14)
where $M$ is the number of preference questions, $\bar{a}^{(i)}$ is the mean of participant $i$’s answers, and $\bar{\bar{a}}$ is the mean of the overall mean vector $\bar{\mathbf{a}}$.
We exclude participant $i$ if $r^{(i)}<0.4$, considering such responses to be poorly correlated with the overall trend. In the end, we retained 22 valid participant responses, and calculated the final Mean Opinion Scores (MOS) based on these.
A.2 Experimental Video Categories
To analyze the impact of different viewpoint conditions on spatial audio perception, we categorized the evaluation videos into two major groups: static and dynamic viewpoints.
A video is considered to have a static viewpoint if the camera position is fixed throughout the clip, or undergoes only minor transformations of original camera pose but still remains fixed. In contrast, a dynamic viewpoint involves explicit camera motion, such as linear translation, arc-shaped trajectories around a subject, or push-in/pull-back zooming effects that create continuous changes in perspective.
To ensure a balanced evaluation, we selected a diverse set of examples from each category, as summarized in Table2. The dataset includes a roughly balanced distribution of static and dynamic scenarios.
View Type Subtype Number of Samples
Static Raw viewpoint (static sound source) 3
Raw viewpoint (dynamic sound source) 7
Fixed camera (transformation) 4
Total (Static) 14
Dynamic Translational movement 5
Arc-shaped trajectory 7
Push-in / Pull-back 7
Total (Dynamic) 19
Table 2: Distribution of evaluation videos by viewpoint category
Appendix B Details of Physics-Driven Spatial Audio Synthesis
Room Impulse Response (RIR) plays a crucial role in spatial audio rendering by modeling how sound propagates and reflects within an environment. Accurate RIR simulation enables the synthesis of physically plausible spatial cues, enhancing the realism of the generated audio experience.
In our pipeline, we adopt gpuRIR [59] as the simulator to generate synthetic RIRs. To reflect varying acoustic environments, we predefine four types of room size categories with randomly sampled dimensions in meters, as shown below:
  • Small: Length and width $\in$ [3, 8], height $\in$ [2.5, 4]
  • Medium: Length and width $\in$ [8, 15], height $\in$ [3, 6]
  • Large: Length and width $\in$ [15, 30], height $\in$ [5, 10]
  • Outdoor: Length and width $\in$ [100, 200], height $\in$ [50, 100]
For each category, the length, width, and height are randomly sampled within the corresponding range. Given the source and receiver trajectories derived in Stage 2, we compute the minimal bounding cuboid that encapsulates both trajectories. This cuboid is then placed at the center of the room. If the bounding box exceeds the selected room size, it is uniformly scaled down to fit the lower bounds of the selected category.
To ensure smooth transitions and minimize auditory artifacts, we perform temporal upsampling on the trajectories before simulation. Specifically, we use a high-resolution sampling rate of $n=44100$ points per trajectory.
For microphone orientation, we support two polar patterns: omnidirectional (omni) and cardioid (card). In the cardioid setting, the microphone orientation at each point is aligned to face outward from the center plane of the rendering space.
Finally, the reverberation time ($T_{60}$) is randomly sampled from a uniform distribution in the range (0.3, 0.6) seconds to simulate varying acoustic characteristics across environments.